A rectangular surface with horizontal dimensions of 4.0 cm and 5.0 cm is found in a uniform magnetic field directed at a 40° angle above the horizontal. Calculate the magnitude of the magnetic field that will yield a flux equivalent to 6.17 ✕ 10 -4 Wb.


1
Expert's answer
2022-04-06T13:59:46-0400

Given:

"A=\rm 4.0\: cm\times 5.0\: cm=0.002\: m^2"

"\theta=90^\circ-40^\circ=50^\circ"

"\Phi=6.17 \times 10^{-4}\:\rm Wb"


The magnitude of magnetic field

"B=\frac{\Phi}{A\cos\theta}"

"B=\frac{6.17 \times 10^{-4}}{0.002*\cos50^\circ}=0.48\:\rm T"